OpenStudy (anonymous):

What is the value of x? 12 - 3/5x - 1/5x = -16

OpenStudy (kamibug):

Combine Like Terms. 12 - 2/5x = -16 Subtract 12 from both sides. -2/5x = -28 Divide both sides by -2/5. x = -28 / -0.4 x = ?

OpenStudy (anonymous):

how did u get 2/5

OpenStudy (anonymous):

i meant 2/5x

OpenStudy (kamibug):

When we combined like terms we subtracted 1/5 from 3/5. :)